# Heuristic search for linear positive systems

### v1

@@ 5 Heuristic search for positive systems
Figure 5: Illustration of the search space after the first (left) and ninth (right) iteration of Algorithm 1 for the system in Example 6.1. States in dark blue are part of the search space $S$ while shaded states are in the active boundary of states that can be directly affected by the dynamics inside $S$. The red circle indicates the state selected for inclusion in the next iteration. The accompanying plots show the optimal cost function $p$ (dashed red) for each state, as well as the current estimates $\overline{g}$ and $\underline{g}$ (yellow) for states in $S$ and the heuristic bounds $\overline{h}$ and $\underline{h}$ (blue).
@@ 6 Numerical example
To give a plausible example of a process that satisfies the necessary assumptions, we consider a large chemical plant. The state $x\in\mathbb{R}^{25}$ represents amounts of $n=25$ different chemical waste compounds that are to be disposed at as low cost (corresponding to energy consumption, pollution or resources consumed) as possible. For each compound a variety of reactions can be induced, represented by the dynamics $Ax+Bu$, which give the resultant compounds as the updated state. We generate a random matrix ${A\geq 0}$ with some sparsity and construct ${B\in\mathbb{R}^{25\times 75}}$ such that ${E=A}$ guarantees positivity of the dynamics, see (Ohlin et al., 2024).
Let $m_{i}=2$ for all $i$ giving a total of three reactions, one represented by the autonomous dynamics and the other two by inputs $u_{ij}$. The cost of each reaction is modeled by vectors $s$ and $r$. Further, the first action for each state $u_{i1}$ is modeled as a direct disposal of all compound of the corresponding state at a high cost. This could represent e.g. burning or releasing the dangerous chemical, resulting in large amounts of pollution. In the input matrix $B$ this is modeled as the column multiplying $u_{i1}$ having its only non-zero element $-1$ on its $i$th row. The initial stabilizing policy $\hat{K}$ is constructed by taking the action $u_{i1}$ for all $i$, giving a suboptimally high cost but guaranteeing that the state is stable. As a consequence, $A+B\hat{K}=\mathbf{0}$. The expression for the upper bound heuristic in Proposition 4.3 then becomes
@@
We note that this is equivalent to $\overline{h}=s+A^{\top}\hat{r}$ where $\hat{r}\in\mathbb{R}^{25}$ is the subvector of entries of $r$ corresponding to the initial policy. As the lower bound policy is given by $\underline{h}=s$, the initial uncertainty in the cost function is quantified by $A^{\top}\hat{r}$.
@@ 7 Conclusion
The presented work builds on recent results for positive linear systems with explicit solutions. This allows us to demonstrate an equivalence between a classes of such problems and the class of stochastic shortest path problems that are the focus of heuristic search methods. An example of such a method applied to the optimal control setting is given, which gives results for the full class of optimal control problems, broadening the scope of applicability beyond SSP. This connection can serve as a bridge to gain a wider field of application for existing results in both areas. Specifically, we argue for the possibility of designing heuristic search algorithms inspired by existing methods for shortest path problems to achieve scalable distributed control of large-scale networked systems.
@@
We would like to thank Olle Kjellqvist and Jonas Hansson at Lund university for insightful discussion related to the presented material.
Appendix A Proof of Theorem 1
Throughout the proof we utilize the fact that the positive orthant $\mathbb{R}^{n}_{\geq 0}$ is invariant under the dynamics as a direct consequence of Assumption 1. Let $J_{k}(x)=p_{k}^{\top}x$. Given the ansatz $p_{0}=0$, we compute $p_{k+1}$ by Value Iteration
$p_{k+1}^{\top}x$ $=\min_{u\in U(x)}\left[g(x,u)+J_{k}(f(x,u))\right]$
$=\min_{u\in U(x)}\left[s^{\top}x+r^{\top}u+p_{k}^{\top}(Ax+Bu)\right]$
$=s^{\top}x+p_{k}^{\top}Ax+\min_{u\in U(x)}\left[(r+B^{\top}p_{k})^{\top}u\right]$
where $U(x)$ denotes the set of all inputs $u$ satisfying the constraints in (1). We partition the vector $r+B^{\top}p_{k}$ in the same fashion as $u$ into subvectors $r_{i}+B_{i}^{\top}p_{k}\in\mathbb{R}^{m_{i}}$. Since each subvector $u_{i}$ is only bound by one of the upper input constraints, we minimize the expression above independently for each $u_{i}$:
$\min_{u\in U(x)}\!(r+B^{\top}p_{k})^{\top}\!u$ $=\!\sum_{i=1}^{M}\min_{\begin{subarray}{c}\mathbf{1}^{\top}u_{i}\leq E_{i}^{\top}x\\ u\geq 0\end{subarray}}\!(r_{i}+B_{i}^{\top}p_{k})^{\top}u_{i}$ (14)
$=\!\sum_{i=1}^{M}\min\{r_{i}+B_{i}^{\top}p_{k},0\}E_{i}^{\top}x$
recalling that the minimum is taken over all elements in the vector $r_{i}+B_{i}^{\top}p_{k}$ and 0. The iteration becomes
$\min_{u\in U(x)}\!\left[g(x,u)+J_{k}(f(x,u))\right]$
$=s^{\top}x+p_{k}^{\top}Ax+\sum_{i=1}^{M}\min\{r_{i}+B_{i}^{\top}p_{k},0\}E_{i}^{\top}x$
$=\left(s+A^{\top}p_{k}+\sum_{i=1}^{M}\min\{(r_{i}+B_{i}^{\top}p_{k},0\}E_{i}\right)^{\top}\!\!x$
$=p_{k+1}^{\top}x.$
Given that (1) has a finite value according to ($i$), the sequence of $J_{k}(x)=p_{k}^{\top}x$ has the upper limit ${J^{*}(x)=p^{\top}x}$. This limit fulfills the equation in ($ii$). The converse also holds since, given a $p$ satisfying ($ii$), $\{p_{k}\}_{k=0}^{\infty}$ must satisfy ${p_{k}\leq p}$ for all $k$. This means that the Bellman equation has a nonnegative and finite solution, implying ($i$). To find the optimal controller we seek the policy $\mu(x)$ minimizing (14) for the optimal cost ${J^{*}(x)=p^{\top}x}$. We get
$\mu(x)=\textnormal{arg}\!\!\!\!\min_{u\in U(x)}\!(r+B^{\top}p)^{\top}u$
which, partitioned like (14), can be performed separately for each subvector $\mu_{i}$ yielding
$\mu_{i}(x)=\underbrace{\begin{bmatrix}\mathbf{0}_{j-1\times n}\\ E_{i}^{\top}\\ \mathbf{0}_{m_{i}-j\times n}\end{bmatrix}}_{=K_{i}}x.$
Here, $E_{i}^{\top}$ enters on the $j$th row and ${j\in\{1,...,m_{i}\}}$ is the index of the largest in magnitude negative element of ${r_{i}+B_{i}^{\top}p}$. If all elements are positive then ${\mu_{i}(x)=\mathbf{0}_{m_{i}\times n}}$. An optimal state feedback law $u_{i}$ is given by ${u_{i}(t)=K_{i}x(t)}$. This shows that the optimal cost is always attained for some ${K\in\mathcal{K}}$, as expected given the linear cost and homogeneous constraints of (1).
Finally, if ($ii$) holds then the $p$ fulfilling (2) clearly also solves the program
Maximize    $\;\;\;\mathbf{1}^{\top}p\;\textnormal{over}\;p\in\mathbb{R}^{n}_{+}$
subject to $\;\;\;p\leq s+A^{\top}p+\sum_{i=1}^{M}\min\{r_{i}+B_{i}^{\top}p,0\}E_{i}$
which is equivalent to the program in ($iii$). Any solution fulfilling the Bellman inequality gives a lower bound on the optimal cost function. Thus the solution to the linear program must also be the optimal cost function and ($ii$) implies ($iii$). To show the converse, assume that the linear program in ($iii$) is bounded by some $p_{k}$. Calculate
$p_{k+1}=s+A^{\top}p_{k}+\sum_{i=1}^{M}\min\{r_{i}+B_{i}^{\top}p_{k},0\}E_{i}.$
We now have $p_{k+1}\geq p_{k}$ and as a consequence
$p_{k+1}$ $=s+A^{\top}p_{k}+\sum_{i=1}^{M}\min\{r_{i}+B_{i}^{\top}p_{k},0\}E_{i}$
$=\min_{u\in U(x)}\left[s^{\top}x+r^{\top}u+p_{k}^{\top}(Ax+Bu)\right]$
$\leq\min_{u\in U(x)}\left[s^{\top}x+r^{\top}u+p_{k+1}^{\top}(Ax+Bu)\right]$
$=s+A^{\top}p_{k+1}+\sum_{i=1}^{M}\min\{r_{i}+B_{i}^{\top}p_{k+1},0\}E_{i}$
which is exactly the constraint in the reformulated program above. Since $p_{k}$ attains the maximum of the program and $p_{k+1}\geq p_{k}$, the only possible conclusion is that ${p_{k+1}=p_{k}}$. Thus the maximizing argument of the program in ($iii$), if it is bounded, solves Equation 2, implying ($ii$). This concludes the proof. $\blacksquare$

### v2

@@ 6 Distributed value iteration
Figure 4: Illustration of the search space after the first (left) and ninth (right) iteration of Algorithm 1 for the system in Example 7.1. Above, states in dark blue are part of the search space $S$ while shaded states are in the active boundary of states that can be directly affected by the dynamics inside $S$. The red circle indicates the state selected for inclusion in the next iteration. The accompanying plots show the optimal cost function $p$ (dashed red) for each state, as well as the current estimates $\overline{g}$ and $\underline{g}$ (yellow) for states in $S$ and the heuristic bounds $\overline{h}$ and $\underline{h}$ (blue). It is clear from the schematic view of the state space that the algorithm systematically explores states most impacted from the initial configuration. Note in particular that the performance of $x_{3}$ is quickly controlled to within the performance bound, while $x_{2}$ remains uncertain, necessitating the exploration of a large part of the remaining state space.
Figure 5: Upper and lower bound of the total cost in Example 7.1. The red lines highlight the iterations for which the state is displayed in more detail in Figure 4. The value $\gamma$ specifies the stopping condition in Algorithm 1 and guarantees a performance of at least $\overline{g}^{\top}x_{0}\leq\gamma p^{\top}x_{0}$ when the algorithm terminates. Since the initial heuristics in Example 7.1 are consistent, the convergence is monotone.
@@ 7 Numerical example
To give a plausible example of a process that satisfies the necessary assumptions, we consider a large chemical plant. The state $x\in\mathbb{R}^{25}$ represents amounts of $n=25$ different chemical waste compounds that are to be disposed at as low cost (corresponding to energy consumption, pollution or resources consumed) as possible. For each compound a variety of reactions can be induced, represented by the dynamics $Ax+Bu$, which give the resultant compounds as the updated state. We generate a random matrix $A\geq 0$} with some sparsity and construct $B\in\mathbb{R}^{25\times 75}$} such that E=A{E=A} guarantees positivity of the dynamics, see (Ohlin et al., 2024).
Let $m_{i}=2$ for all $i$ giving a total of three reactions, one represented by the autonomous dynamics and the other two by inputs $u_{ij}$. The cost of each reaction is modeled by vectors $s$ and $r$. Further, the first action for each state $u_{i1}$ is modeled as a direct disposal of all compound of the corresponding state at a high cost. This could represent e.g. burning or releasing the dangerous chemical, resulting in large amounts of pollution. In the input matrix $B$ this is modeled as the column multiplying $u_{i1}$ having its only non-zero element 1-1 on its $i$th row. The initial stabilizing policy $\hat{K}$ is constructed by taking the action $u_{i1}$ for all $i$, giving a suboptimally high cost but guaranteeing that the state is stable. As a consequence, $A+B\hat{K}=\mathbf{0}$. The expression for the upper bound heuristic in Proposition 4.2 then becomes
@@
Note that, due to the above construction of $B$, this is equivalent to $\overline{h}=s+A^{\top}\hat{r}$ where $\hat{r}\in\mathbb{R}^{25}$ is the subvector of entries of $r$ corresponding to the initial policy. As the lower bound policy is given by $\underline{h}=s$, the initial uncertainty in the cost function is quantified by $A^{\top}\hat{r}$.
@@ 8 Conclusion
The presented work builds on recent results for positive linear systems with explicit solutions. This allows us to demonstrate an equivalence between a subset of such problems and the class of stochastic shortest path problems that are the focus of heuristic search methods. An example of such a method applied to the optimal control setting is given, which gives results for the full class of optimal control problems, broadening the scope of applicability beyond SSP. This connection can serve as a bridge to gain a wider field of application for existing results in both areas. An algorithm for distributed optimal control is presented, where the derived heuristic bounds are the key tool to enable global bounds on performance.
We argue for the possibility of designing heuristic search algorithms, with several initial examples provided here, inspired by existing methods for shortest path problems to achieve scalable distributed control of large-scale networked systems. Promising areas of further study include the dual to the presented problem, which suggests that favorable results could be derived for state observers of positive systems.
@@
We would like to thank Olle Kjellqvist, Jonas Hansson and Mark Jeeninga at Lund university and César Uribe at Rice Academy for insightful discussion related to the presented material. We are also grateful to the anonymous reviewers for their constructive comments and helpful advice.